OpenStudy (anonymous):

Solve this system. Please put your answer as an ordered pair. -x - 7y = -14 -4x - 14y = -28

OpenStudy (anonymous):

Well

OpenStudy (anonymous):

x=0. y=2.

OpenStudy (anonymous):

sorry i have to go

OpenStudy (anonymous):

ill be back i have to help some peeps

OpenStudy (bradely):

multiply the first equation by (4) -4x-28y=-56 -----(1) -4x-14y=-56 -----(2) -------- (1)-(2) -14y=0 y=0 -4x=-56 x=14 If any doubts contact me www.mathskey.com